OpenStudy (anonymous):

I WILL GIVE A MEDAL OF WHO CAN HELP ME SOLVE THIS EQUATION(it is in the comments)

OpenStudy (anonymous):

What is the area of the figure?

OpenStudy (anonymous):

@ankit042

OpenStudy (ankit042):

Can you identify different rectangles/squares in the figure?

OpenStudy (anonymous):

yes

OpenStudy (ankit042):

Great can you tell all the figures (hint there are 3)

OpenStudy (anonymous):

3 rectangles? or 2 rectanges and square

OpenStudy (anonymous):

just 3 rectangles

OpenStudy (ankit042):

Yes there are 3 rectangles now area for rectangle is length*breath

OpenStudy (ankit042):

so 5*12+5*12+5*4

OpenStudy (anonymous):

Is the answer 140?

OpenStudy (ankit042):

Yes!!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Which phrase best defines a parallelogram? A. four congruent sides B. at least two right angles C. four congruent angles D. two pairs of parallel sides

OpenStudy (anonymous):

@ankit042 @Mertsj @HelpBlahBlahBlah

OpenStudy (ankit042):

This is simple definition why don't you give it a try!!

OpenStudy (anonymous):

I am really bad at math

OpenStudy (anonymous):

d?

OpenStudy (ankit042):

LOL this is not Math just english :)

OpenStudy (helpblahblahblah):

i think its D

OpenStudy (anonymous):

A parallelogram is a simple quadrilateral with two pairs of parallel sides :) says wiki.

OpenStudy (ankit042):

yes see you answered it correctly :D

OpenStudy (anonymous):

thxkamizamurai and ankit and help blahblahblah :D

OpenStudy (helpblahblahblah):

anytime :)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

@ankit042 @mathmale @HelpBlahBlahBlah @Kamizamurai Which equation is true? A. 529 square root=529/2 B.529 square root =23 squared square root C.529 square root=529*2 D.529 square root=23 squared

OpenStudy (ankit042):

Do you know the definition of square root?

OpenStudy (anonymous):

@LeoRay

OpenStudy (anonymous):

Sort of it means multipliing by 3

OpenStudy (anonymous):

right?

OpenStudy (ankit042):

NOooooo

OpenStudy (anonymous):

x * x = 529 x = ?

OpenStudy (ankit042):

Yes @Kamizamurai did it correctly!

OpenStudy (anonymous):

264.5

OpenStudy (anonymous):

264.5 * 264.5 = 69960.25

OpenStudy (anonymous):

ok 1 second

OpenStudy (anonymous):

ummmmmm how i am SOOOO confused

OpenStudy (leoray):

23 * 23 is 529

OpenStudy (anonymous):

OH 23? ok

OpenStudy (leoray):

Here. You can learn more through this link : http://www.homeschoolmath.net/teaching/square-root-algorithm.php

OpenStudy (leoray):

The answer is D. The square root of 529 is 23.

OpenStudy (anonymous):

What heck? i posted thhat question a while ago and it shows up now? i get it though

OpenStudy (leoray):

You're welcome.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Can someone give Leoray a medel?

OpenStudy (anonymous):

What is the area of this rectangle in square centimeters? A. 3600 square centimeters B. 36 square centimeters C. 360,000 square centimeters D. 36,000 square centimeters

OpenStudy (leoray):

The answer is B. 9*4=36 A= W*L

OpenStudy (anonymous):

thx

OpenStudy (leoray):

You're welcome.